# Machine Unlearning: A Comprehensive Survey

### v1

@@ 2.2. Unlearning Evaluation Metrics
To compare the two models before and after unlearning, we need to define an evaluating metric $d(.)$ between $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ from the model parameter space $\mathcal{H}$. To this end, we briefly introduce several common evaluation distances metrics:
@@
$L_{2}$-Norm. Wu et al. (Wu et al., 2020) proposed to utilise the Euclidean norm to evaluate the parameters of $\mathcal{A}(D\backslash D_{e})$ and that of unlearning $\mathcal{U}(D,D_{e},\mathcal{A}(D))$. Suppose $\theta$ are the model parameters learned by algorithm $\mathcal{A}(.)$. The $l_{2}$-norm or distance can be denoted as $||\theta_{\mathcal{A}(D\backslash D_{e})}-\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}||$, where $\theta_{\mathcal{A}(D\backslash D_{e})}$ is the model parameters of retraining from the scratch and $\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}$ is the modal parameters of unlearning algorithms $\mathcal{U}(.)$.
@@
Kullback–Leibler divergence (KLD). In Bayes-based or Markov chain Monte Carlo-based unlearning methods, researchers used the Kullback–Leibler divergence or Jensen-Shannon divergence to optimize the approximate posterior since they can measure the distance of two probability distributions. Therefore, in recent unlearning studies, researchers also used the KLD between the retrained and unlearned models to estimate the unlearning effect (Nguyen et al., 2020). For discrete probability distritbutions of two model parameters $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ on the same probability space $\mathcal{H}$, the KLD from $\mathcal{A}(D\backslash D_{e})$ to $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is denoted as $\text{KL}[\Pr(\mathcal{A}(D\backslash D_{e}))||\Pr(\mathcal{U}(D,D_{e},\mathcal{A}(D)))]$.
@@
Parameter Leakage. Since membership inference attacks can infer if a sample is used in the training dataset of a model, recently, some works have leveraged the property to verify if the specific data is removed by unlearning mechanisms. Some studies (Hu et al., 2022) even proposed to backdoor the erased samples and then attack the unlearned model to check if the unlearned model can be attacked. If the unlearned model is backdoored, it proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022; Ma et al., 2022) to evaluate the unlearning effect.
@@ 2.3. Majority of Tools Used in Unlearning
2.3.1. Differential Privacy (DP)
Differential privacy has become a popular standard for privacy protection in the statistic. In a DP model, a trusted analyzer collects users’ raw data and then executes a private mechanism to achieve differential privacy. The privacy requirement ensures the indistinguishability for any two outputs of neighbouring datasets, where neighbouring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$. A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every pair of neighboring datasets $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate. The smaller the $\epsilon$ is, the more the privacy protection level increases. When the value of $\epsilon$ is 0, it implies that the outputted probability distribution of algorithm $\mathcal{M}$ cannot reflect any useful information about the datasets. A general DP mechanism based on adding Laplace noise was presented and theoretically analyzed in (Dwork et al., 2006).
2.3.2. Bayesian Variational Inference (VI)
In machine learning, the bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization. In this part, we revisit the VI framework (Blei et al., 2017) that learns approximate posterior model parameters $\theta$ using Bayesian Theory (Bernardo and Smith, 2009). Given a prior belief $p(\theta)$ of an unknown model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the Kullback–Leibler divergence (KLD) (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$. The KLD is a type of statistical distance to measure how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$. However, it is intractable to compute the KLD exactly or minimize the KLD directly. Instead, the evidence lower bound (ELBO) (Blei et al., 2017) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions. ELBO follows directly from $log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where the log marginal likelihood $log(p(D))$ is independent of $q(\theta|D)$. Since $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$, the ELBO is a lower bound of $log(p(D))$.
In general training situations, ELBO is maximized using stochastic gradient ascent (SGA). The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[log(p(D|\theta))+log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA. We can use a simple distribution (e.g., the exponential family)to approximate computational ease posterior belief $q(\theta|D)$. When dealing with complex distribution whose density cannot be evaluated, we can maximize the ELBO with adversarial training by alternating between estimating the log-density ratio $log(p(\theta))/q(\theta|D)$ and maximizing the ELBO (Yu et al., 2019).
2.3.3. Privacy Leakage
@@
Privacy leakage occurs in both unlearning verification and privacy threats in unlearning two parts. In unlearning verification, researchers tried to use privacy leakage attacks to verify if the specific data is unlearned. In the privacy and security issues in unlearning, researchers tried to design effective inference attacks tailored to machine unlearning. The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a data sample was used in the ML model updating process or not. When a sample is fully known to the adversary, learning which model was trained based on it will leak information about the model. A generic membership inference process was introduced in (Shokri et al., 2017). Shokri et al. first trained the shadow models to simulate the target machine-learning model. Then, they observed and stored the different outputs of the shadow models based on different inputs, in or not, in the trainset. They used these stored outputs as samples to train the membership inference attack model.
@@
Model inversion (Fredrikson et al., 2014), or privacy reconstruction (Salem et al., 2020) is another privacy threat in general machine learning. The goal of model inversion is to infer some missing attributes of input features based on the interaction with a trained machine learning model. Salem et al. (Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating process by the different outputs of the model before updating and after updating. Latter, inferring the private information of updating data in conventional machine learning is transferred to infer the privacy of the erased samples in machine unlearning. In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{probe}$, including the original outputs $\hat{Y}_{\mathcal{M}}$ before unlearning, and the outputs $\hat{Y}_{\mathcal{M}_{-D_{e}}}$ after unlearning. Then he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{\mathcal{M}_{-D_{e}}}-\hat{Y}_{\mathcal{M}}$. The attack model contains an encoder and decoder, which has a similar structure as VAEs (Kingma and Welling, 2014), and the main process is shown in Fig. 3.
3. Categories of Centralized Machine Unlearning
@@
In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes. Detailed techniques will be introduced in the next section. From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it needs to store the whole training dataset and retrain a new model from scratch. It consumes a huge amount of storage and computation costs, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms. However, there were difficult challenges when we tried other unlearning methods. Three representative challenges are the stochasticity of training, the incrementality of training, and the catastrophe of unlearning. Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets. It also divides the stochasticity and the incrementality into sub-models. It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining. Another approach is approximate unlearning. It can reduce both storage and computation costs because it unlearns only based on the erased data. However, it is difficult for approximate unlearning methods to control the accuracy degradation, due to the challenges in estimating stochasticity and incrementality during the training process. Most of them bounded their estimation to avoid dramatic accuracy reduction.
3.1. Exact Unlearning
@@
Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background introduction of unlearning, we know the learning model $\mathcal{A}$ and unlearning model $\mathcal{U}$ based on the trainset $D$ and erased dataset $D_{e}\subseteq D$. If $\mathcal{U}(.)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\backslash D_{e})\in\mathcal{H}$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022). Unlike naive retraining needs to retrain based on the whole remaining dataset, exact unlearning tries to retrain a sub-model based on the subset of remaining data to reduce computation cost. A general operation of exact unlearning is that they first divide the training dataset into several sub-sets. Then, they transform the learning process by ensembling the sub-models trained based on each sub-set as the final model. So that when unlearning request comes, they only need to retrain the trained sub-model based on the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
@@
As we introduced above, we know that the primary goal of exact unlearning is to reduce the computation cost of retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form. In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. When an unlearning requirement comes, they only need to update several summations, ensuring the process runs faster than retraining from scratch. Bourtoule et al. proposed SISA (Bourtoule et al., 2021), which is short for Sharded, Isolated, Sliced, and Aggregated training. SISA divided the full data into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data. Wu et al. (Wu et al., 2022a) proposed a framework that precisely models the influence of each training data point on the model concerning various performance criteria and removes the influence of the data that is required to delete. Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not delete, and the unlearning part on the erased data with parameters bound. These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
@@
Besides the huge storage cost, another big problem in exact unlearning is that it only suits the situation, in which the unlearning request only needs to remove a few samples with a low frequency. Suppose an unlearning request needs to remove many data samples (usually, they are not in the same previous divided sub-set). In that case, exact unlearning must retrains all these related sub-models or even all the sub-models in the worst situation. At the moment, exact unlearning is no longer computation efficient, and it still needs to store the whole training dataset or intermediate parameters.
3.2. Approximate Unlearning
@@
Figure 4. Model changes when adding a new point or removing a point. (a) A normally trained classifying model classifies classes 1 and 2. (b) When a new point appears, the model is trained based on it, and the classifying line is pushed to classify it. (c) When we need to remove an influential point, we should recover the contribution of this point to the model. (d) When we remove a Non-influential point, the model may not need to change a lot.
Unlike exact unlearning could only reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the learned model and the erased data sample, which saves the computation and storage costs together. The goal of approximate unlearning is to unlearn a posterior that approximates the model trained on the remaining dataset, based on only the erased sub-dataset while not the remaining dataset (used by exact unlearning).
@@
As we introduced before, the unlearning purpose is expecting that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ to equal to the retrained model $\mathcal{A}(D\backslash D_{e})$. Since exact unlearning is implemented by retraining from the remaining dataset or sub-sets, they can almost guarantee equality before and after unlearning. However, since approximate unlearning tries to directly delete the contribution of the erased data from the trained model, how to exactly estimate the contribution of the erased data (the contribution contains the stochasticity and the incrementality ) and remove it becomes the core problem.
@@
The text description of the changes between two different distribution spaces before and after removing the specific data is not intuitive. Fig. 4 shows illustrated changes when adding a new point or removing a point in a classifying model. When an influential point appears, it usually pushes the line to move forward than the original classifying line to identify it, as shown in Fig. 4 (b). When this influential point is requested to remove, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig. 4 (c). However, when only unlearning a non-influential point, which may have almost non-influence on the model, the unlearned model may not change compared to the original trained model in this situation, as shown in Fig. 4 (d).
@@
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal and Bayes-based mechanisms, which will be introduced in technical detail in Section 4. Here, we briefly summarize current studies about approximate unlearning. A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy. Guo et al. (Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP. The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning. The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021). Sekhari et al. (Sekhari et al., 2021) introduced a similar solution to (Guo et al., 2020), which removes the contribution of the erased data from the gradients during the unlearning process. To mitigate the unlearning accuracy reduction, they bounded the model parameters by their differential unlearning definition. Neel et al. (Neel et al., 2021) proposed unlearning methods allowing arbitrary sequences of updates, including addition and deletion, and defined the corresponding weak and strong unlearning algorithms.
@@
Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior directly. In (Nguyen et al., 2020), Nguyen et al. unlearned an approximate posterior based on the erased data via the Bayesian inference. Similarly, Fu et al. (Fu et al., 2022) proposed a machine unlearning algorithm for Markov chain Monte Carlo (MCMC) and further defined a knowledge removal estimator the evaluate the knowledge erasing. Nguyen et al. (Nguyen et al., 2022a) also explained the effectiveness of the approximate machine unlearning algorithm from the perspective of Markov Chain Monte Carlo. However, since those methods are approximately unlearning the contribution of all the erasing data, including the inputs and labels, it is easy to decrease the model accuracy to some extent after unlearning.
4. Techniques of Centralized Unlearning
Table 1. Summary of popular unlearning techniques

### v2

@@ 3.2. Machine Unlearning Evaluation Metrics
To compare the two models before and after unlearning, we need to define an evaluating metric $d(\cdot)$ between $\mathcal{A}(D\setminus D_{e})$ and $\mathcal{U}(M,D,D_{e})$. To this end, we briefly introduce several common evaluation distance metrics:
@@
$L_{2}$-Norm. In (Wu et al., 2020), the authors propose utilizing the Euclidean distance to evaluate the parameters of the retrained model and the unlearned model. Let $\theta$ represent the model parameters learned by the algorithm $\mathcal{A}(\cdot)$. The $L_{2}$-norm measures the distance between $\theta_{\mathcal{A}(D\setminus D_{e})}$ and $\theta_{\mathcal{U}(M,D,D_{e})}$, where $\theta_{\mathcal{A}(D\setminus D_{e})}$ are the model parameters retrained from scratch, and $\theta_{\mathcal{U}(M,D,D_{e})}$ are the model parameters resulting from the unlearning algorithm $\mathcal{U}(\cdot)$.
@@
Kullback–Leibler divergence (KLD). KLD is commonly used to measure the divergence between two probability distributions, often assessing the distance between retrained and unlearned models. In Bayes-based or Markov chain Monte Carlo-based unlearning methods (Nguyen et al., 2020), researchers utilize KLD (Joyce, 2011) to optimize approximate models, employing it to measure the distance between two probability distributions. Recent unlearning studies have also used KLD to estimate the unlearning effect by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
@@
Evaluation Metric based on Privacy Leakage. Since membership inference attacks (Chen et al., 2021) can decide whether a sample was utilized for training a model, recently, some works have leveraged this property to verify if unlearning mechanisms remove the specific data. Some studies (Hu et al., 2022) even proposed to backdoor the unlearning samples for initial model training and then attack the unlearned model. If the unlearned model is still backdoored, this proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effect.
@@ 3.3. Majority of Tools Used in Unlearning
Differential Privacy (DP). Differential privacy is a popular benchmark for privacy protection in the Statistic (Dwork et al., 2006). In a DP model, a trusted analyzer collects users’ raw data and then executes a private method to guarantee differential privacy. The DP protection ensures the indistinguishability for any two outputs of neighboring datasets, where neighboring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$. A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every neighboring dataset pair $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate. The degree of privacy protection rises with decreasing $\epsilon$. When $\epsilon=0$, it implies that the outputted probability distribution of mechanism $\mathcal{M}$ cannot represent any meaningful information. A general DP mechanism based on adding Laplace noise was presented and theoretically analyzed in (Dwork et al., 2006).
Bayesian Variational Inference. In machine learning, the Bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization (Kingma and Welling, 2014; Blei et al., 2017). We revisit the variational inference framework that learns approximate posterior model parameters $\theta$ using Bayesian Theory in this part. Suppose a prior belief $p(\theta)$ of an unidentified model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the KLD (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$. KLD measures how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$. However, it is intractable to compute the KLD exactly or minimize the KLD directly. Instead, the evidence lower bound (ELBO) (Kingma and Welling, 2014) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions. ELBO follows directly from $\log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where $\log(p(D))$ is independent of $q(\theta|D)$. The ELBO is a lower bound of $\log(p(D))$ as $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$. In general training situations, ELBO is maximized using stochastic gradient ascent (SGA) (Kingma and Welling, 2014). The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[\log(p(D|\theta))+\log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA. We can use a simple distribution (e.g., the exponential family)to approximate computational ease posterior belief $q(\theta|D)$.
@@
Privacy Leakage Attacks. Privacy leakage occurs in both unlearning verification and privacy threats in two parts of unlearning. In unlearning verification, researchers tried to use privacy leakage attacks to verify whether the specific data is unlearned. Regarding the privacy and security issues in unlearning, researchers have tried to design effective inference attacks tailored to machine unlearning. The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a sample was employed in the model updating process or not. When an attacker fully knows a sample, knowing which model was trained on it will leak information about the model. A generic membership inference process was introduced in (Shokri et al., 2017). Shokri et al. first trained the shadow models to approach the target ML models. Then, they observed and stored the different outputs of the shadow models based on different inputs, in or not, in the trainset. They used these stored outputs as samples to train the membership inference attack model.
@@
Model inversion (Fredrikson et al., 2014), or privacy reconstruction (Salem et al., 2020) is another privacy threat in general machine learning. Model inversion aims to infer some lacking attributes of input features based on the interaction with the trained ML model. Salem et al. (Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating by different model outputs before and after updating. Later, inferring the private information of updating data in conventional machine learning is transferred to inferring the privacy of the erased samples in machine unlearning. In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{\text{probe}}$, including the original outputs $\hat{Y}_{M}$ before unlearning, and the outputs $\hat{Y}_{M_{-D_{e}}}$ after unlearning. Then, he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{M_{-D_{e}}}-\hat{Y}_{M}$. The attack model contains an encoder and decoder, which has a similar structure as VAEs (Kingma and Welling, 2014), and the main process is shown in Fig. 3.
4. Centralized Machine Unlearning
In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
4.1. Unlearning Solution Categories
@@
From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it requires storing the entire original dataset and retraining the model from scratch, which consumes significant storage and computational resources, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms. Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets (Bourtoule et al., 2021). It also divides the stochasticity and the incrementality into sub-models. It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining. Another approach is approximate unlearning (Sekhari et al., 2021). It can reduce both storage and computation costs because it unlearns only based on the erased data. However, it is difficult for approximate unlearning methods to control the accuracy degradation due to the challenges in estimating stochasticity and incrementality during the training process. Most of them bounded their estimation to avoid dramatic accuracy reduction.
4.1.1. Exact Unlearning
@@
Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background introduction of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively. If $\mathcal{U}(\cdot)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\setminus D_{e})\in\mathcal{H}$ and $\mathcal{U}(M,D,D_{e})\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022a). Unlike naive retraining, which relies on the whole remaining dataset, exact unlearning tries to retrain a sub-model only using a subset of the remaining dataset to reduce calculation cost. A general operation of exact unlearning is that they first divide the dataset into several small sub-sets. Then, they transform the learning process by ensembling the sub-models trained with each sub-set as the final model (Bourtoule et al., 2021; Cao and Yang, 2015). So that when an unlearning request comes, they are just required to retrain the sub-model corresponding to the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
@@
As we introduced above, exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form. In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. They are only required to update several summations when an unlearning requirement comes, ensuring the method runs faster than retraining from scratch. SISA (Bourtoule et al., 2021) is a representative exact unlearning algorithm, which splits the full training dataset into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data. Study (Wu et al., 2022a) proposed a framework that precisely models the impact of individual training sample on the model concerning various performance criteria and removes the impact of samples that are required to be removed. Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not be deleted, and the unlearning part with the erased data will be unlearned with parameters bound. These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
@@
Besides the high storage cost, another major issue with exact unlearning is that it is only suitable for scenarios where the unlearning request involves removing a few samples with low frequency. Suppose an unlearning request needs to remove many data samples (usually, they are not in the same previous divided sub-set). In that case, exact unlearning must retrain all these related sub-models or even all the sub-models in the worst situation. At the moment, exact unlearning is no longer computation efficient, and the whole training dataset and intermediate parameters still need to be stored.
4.1.2. Approximate Unlearning
@@
Figure 4. The model changes when adding a new point or removing a point. (a) A normally trained classifying model classifies classes 1 and 2. (b) When a new point appears, the model is trained based on it, and the classifying line is pushed to classify it. (c) When we need to remove an influential point, we should recover the contribution of this data point on the model. (d) When we remove a Non-influential point, the model may not need to change a lot.
@@
Unlike exact unlearning, which only aims to reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the trained model and the erased data sample, which saves the computation and storage costs together. Approximate unlearning studies aim to unlearn a model that approaches the model trained on the remaining dataset, i.e., the unlearned model $\mathcal{U}(M,D,D_{e})$ should match the retrained model $\mathcal{A}(D\setminus D_{e})$. Since exact unlearning is implemented by retraining from the remaining dataset or sub-sets, they can almost guarantee equality before and after unlearning. However, since approximate unlearning tries to directly delete the influence of the unlearned samples from trained models, the core problem lies in precisely estimating and removing this contribution, which includes both stochasticity and incrementality.
@@
The text description of the changes between two different distribution spaces before and after removing the specific data is not intuitive. Fig. 4 shows illustrated changes when adding a new point or removing a point in a classifying model. When an influential point appears, it usually pushes the line to move forward than the original classifying line to identify it, as shown in Fig. 4 (b). When this influential point is requested to be removed, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig. 4 (c). However, when only unlearning a non-influential point, which may have almost non-influence on the model, the unlearned model may not change compared to the original trained model in this situation, as shown in Fig. 4 (d).
@@
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal (Guo et al., 2020) and Bayes-based mechanisms (Nguyen et al., 2020), which are introduced in technical detail in Section 4.2. Here, we briefly summarize current studies about approximate unlearning. A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy. Guo et al. (Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP. The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning. The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021). A similar solution to (Guo et al., 2020) is introduced in (Sekhari et al., 2021), which erases the influence of the specified samples from the gradients during unlearning. To mitigate accuracy degradation, model parameter updates are constrained by a differential unlearning definition.
@@
Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior. In (Nguyen et al., 2020), Nguyen et al. employed Bayesian inference to unlearn an approximate model using the erased data. Moreover, Fu et al. (Fu et al., 2022) developed a similar unlearning method, which uses Markov chain Monte Carlo (MCMC) methods. In (Nguyen et al., 2022a), the authors also explained the effectiveness of the approximate unlearning method from the perspective of MCMC. Nevertheless, since those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
Main Challenges of Approximate Unlearning. As we discussed at the beginning, there are three main challenges in machine unlearning. In centralized scenarios, these studies mainly aim at solving the basic machine unlearning problem, which unavoidably faces three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning. The exact unlearning methods extend the retraining idea, which avoids facing these challenges but consumes lots of storage costs. The approximate unlearning methods face these challenges directly, and we here list the relevant work about how to estimate the contribution of erased samples to overcome the stochasticity and incrementality of training, and how to prevent unlearning catastrophe.
  • To overcome the stochasticity and incrementality challenges when estimating the unlearning influence, one popular strategy is based on the first-order and second-order influence function (Basu et al., 2020), which is calculated based on the perturbation theory (Avrachenkov et al., 2013).
  • The unlearning catastrophe appears commonly in approximate unlearning, and many studies try to propose some methods to solve this problem. In certified removal and Bayesian-based methods, they usually set a threshold to limit the unlearning update extent (Guo et al., 2020; Nguyen et al., 2020). In (Wang et al., 2023a), Wang et al. solves this problem by adding a model utility compensation task during unlearning optimization and finding the optimal balance based on multi-objective training methods.
4.2. Detailed Techniques of Centralized Unlearning
Table 2. Popular Unlearning Techniques